# Metric Distortion of Small-group Deliberation [^1] All Python code is available at http://bit.ly/3WHW8H4

### v3

@@ 2 Model and Preliminaries
We now formally present the model and framework for small group interaction. There is a set $C$ of $m$ candidates (or alternatives or outcomes), and $m!$ possible rankings of these alternatives. We assume a continuum of voters, each of whose preference follows one of the rankings. [^5] We make the continuum assumption on voters for analytic convenience, noting that the relevant lower bounds for the finite voter metric distortion problem hold in the continuum voter setting as well. (See Theorem 2.3.) The preferences of a fraction $\rho_{i}$ of the voters follow ranking $i$.
In the metric distortion framework, we assume the rankings with $\rho_{i}>0$ and the candidates are embedded in a latent metric space with distance function $d(\cdot,\cdot)$. We assume the metric space has discrete (and finite number of) locations, corresponding to the candidates and the rankings with $\rho_{i}>0$. The voters whose preferences follow ranking $i$ are placed at the location corresponding to $i$ [^6] This is without loss of generality – if there are multiple locations for the same ranking, the distortion as defined later weakly increases if these voters are moved to one of the locations. Therefore, each ranking location $i$ is associated with a mass $\rho_{i}>0$ of voters. We use “metric space” to mean both the distance function, as well as the voter mass at each location. We denote the combination $(d,\vec{\rho})$ as $\sigma$.
@@
In the absence of deliberation, we assume the metric space is “consistent” with the rankings. By this, we mean that given location $i$ with $\rho_{i}>0$, and two candidates $c_{1}$ and $c_{2}$, the ranking $i$ places $c_{1}$ higher than $c_{2}$ if $d(i,c_{1})<d(i,c_{2})$. If $d(i,c_{1})=d(i,c_{2})$, then either of the two candidates can be ranked higher than the other. Note that a consistent metric always exists by placing all rankings with $\rho_{i}>0$ at location $a$ and all candidates at location $b$.
A social choice rule $\mathcal{S}$ takes the rankings with $\rho_{i}>0$ (and the corresponding $\rho_{i}$) as input, and outputs a single candidate as the “winner”. Note that the social choice rule only sees the rankings and the $\rho_{i}$, but not the underlying metric space. We also assume the social choice rule is anonymous meaning that its output does not depend on the identities of the alternatives – if the alternatives are permuted by $\pi$, the social choice rule outputs $\pi(c)$ if it was originally outputting $c$.
To measure the quality of the social choice rule, for any candidate $c$, let $SC(c,\sigma)=\sum_{i}\rho_{i}d(i,c)$ denote the social cost of $c$ under the metric $\sigma=(d,\vec{\rho})$. Let $c^{*}(\sigma)=\mbox{argmin}_{c\in C}SC(c,\sigma)$ be the social optimum or the median outcome had the metric space $\sigma$ been known. This outcome need not be unique, in which case an arbitrary such outcome is output. Let $\mathbf{Alg}(\mathcal{S},\vec{\rho})$ denote the output of the social choice rule $\mathcal{S}$ given the rankings $i$ and their masses $\rho_{i}$. Then the distortion of the rule is
@@ 2.1 Deliberation Models
In our deliberative framework, there is a group size parameter $k\geq 2$. A single deliberation is a function that operates over two alternatives $c_{1}$ and $c_{2}$, and $k$ randomly chosen voters from the continuum of voters, and outputs one of the alternatives. By “randomly chosen”, we mean $k$ rankings are chosen independently with replacement from the set of rankings, with ranking $i$ being chosen at each step with probability $\rho_{i}$. Given the continuum assumption on voters, these $k$ rankings (some of which are possibly repeated) correspond to $k$ distinct voters.
In order to define the deliberation function, we first need to define normalized bias. Given a voter at location $i$, define the normalized bias as

### v4

@@ 2 Model and Preliminaries
We now formally present the model and framework for small group interaction. There is a set $C$ of $m$ candidates (or alternatives or outcomes), and $m!$ possible rankings of these alternatives. We assume a continuum of voters, each of whose preference follows one of the rankings. [^5] We make the continuum assumption on voters for analytic convenience, noting that the relevant lower bounds for the finite voter metric distortion problem hold in the continuum voter setting as well. (See Theorem 2.3.) The preferences of a fraction $\rho_{q}$ of the voters follow ranking $q$.
In the metric distortion framework, we assume the rankings with $\rho_{q}>0$ and the candidates are embedded in a latent metric space with distance function $d(\cdot,\cdot)$. We assume the metric space has discrete (and finite number of) locations, corresponding to the candidates and the rankings with $\rho_{q}>0$. The voters whose preferences follow ranking $q$ are placed at a set of locations $S_{q}$, where the voter mass at location $i\in S_{q}$ is $\rho_{iq}$, with $\sum_{i\in S_{q}}\rho_{iq}=\rho_{q}$. We use “metric space” to mean both the distance function, as well as the voter mass at each location. We denote the combination $(d,\vec{\rho})$ as $\sigma$. Let $Q$ denote the set of all locations of voters. Given a location $i\in Q$, the ranking corresponding to $i$ is denoted $\sigma_{i}$, and the voter mass there is denoted $\rho_{i}$.
@@
In the absence of deliberation, we assume the metric space is “consistent” with the rankings. By this, we mean that given location $i$ with $\rho_{i}>0$, and two candidates $c_{1}$ and $c_{2}$, the ranking $q=\sigma_{i}$ places $c_{1}$ higher than $c_{2}$ if $d(i,c_{1})<d(i,c_{2})$. If $d(i,c_{1})=d(i,c_{2})$, then either of the two candidates can be ranked higher than the other. Note that a consistent metric always exists by placing all rankings with $\rho_{q}>0$ at location $a$ and all candidates at location $b$.
A social choice rule $\mathcal{S}$ takes the rankings with $\rho_{q}>0$ (and the corresponding $\rho_{q}$) as input, and outputs a single candidate as the “winner”. Note that the social choice rule only sees the rankings and the $\rho_{q}$, but not the underlying metric space. We also assume the social choice rule is anonymous meaning that its output does not depend on the identities of the alternatives – if the alternatives are permuted by $\pi$, the social choice rule outputs $\pi(c)$ if it was originally outputting $c$.
To measure the quality of the social choice rule, for any candidate $c$, let $SC(c,\sigma)=\sum_{i\in Q}\rho_{i}d(i,c)$ denote the social cost of $c$ under the metric $\sigma=(d,\vec{\rho})$. Let $c^{*}(\sigma)=\mbox{argmin}_{c\in C}SC(c,\sigma)$ be the social optimum or the median outcome had the metric space $\sigma$ been known. This outcome need not be unique, in which case an arbitrary such outcome is output. Let $\mathbf{Alg}(\mathcal{S},\vec{\rho})$ denote the output of the social choice rule $\mathcal{S}$ given the rankings $q$ and their masses $\rho_{q}$. Then the distortion of the rule is
@@ 2.1 Deliberation Models
In our deliberative framework, there is a group size parameter $k\geq 2$. A single deliberation is a function that operates over two alternatives $c_{1}$ and $c_{2}$, and $k$ randomly chosen voters from the continuum of voters, and outputs one of the alternatives. By “randomly chosen”, we mean $k$ rankings are chosen independently with replacement from the set of rankings, with ranking $q$ being chosen at each step with probability $\rho_{q}$. Given the continuum assumption on voters, these $k$ rankings (some of which are possibly repeated) correspond to $k$ distinct voters.
In order to define the deliberation function, we first need to define normalized bias. Given a voter at location $i\in Q$, define the normalized bias as